Costs per Equivalent Unit

The following information concerns production in the Baking Department for March. All direct materials are placed in process at the beginning of production.

ACCOUNT Work in Process—Baking Department ACCOUNT NO.
Date Item Debit Credit Balance
Debit Credit
Mar. 1 Bal., 9,300 units, 4/5 completed 16,182
31 Direct materials, 167,400 units 217,620 233,802
31 Direct labor 63,890 297,692
31 Factory overhead 35,938 333,630
31 Goods finished, 169,500 units 321,678 11,952
31 Bal. ? units, 3/5 completed 11,952

a. Based on the above data, determine each cost listed below. Round "cost per equivalent unit" answers to the nearest cent.

1. Direct materials cost per equivalent unit $
2. Conversion cost per equivalent unit $
3. Cost of the beginning work in process completed during March $
4. Cost of units started and completed during March $
5. Cost of the ending work in process $

b. Assuming that the direct materials cost is the same for February and March, did the conversion cost per equivalent unit increase, decrease, or remain the same in March?
Answer #1
Units charged to production:
Inventory in process, March 1 9300
Received from materials storeroom 167400
Total units accounted for by the Baking Department 176700
Units to be assigned costs:
Equivalent Units
Whole Units Direct Materials Conversion
Inventory in process, March 1 9300 0 1860
Started and completed in March 160200 160200 160200
Transferred to Packing Department in March 169500 160200 162060
Inventory in process, March 31 7200 7200 4320
Total units to be assigned costs 176700 167400 166380
Cost Information
Costs per equivalent unit:
Direct Materials Conversion
Total costs for March in Baking Department 217620 99828
Total equivalent units 167400 166380
Cost per equivalent unit 1.30 0.60
Costs charged to production:
Direct Materials Conversion Total
Inventory in process, March 1 16182
Costs incurred in March 317448
Total costs accounted for by the Baking Department 333630
Cost allocated to completed and partially completed units:
Inventory in process, March 1 balance 16182
To complete inventory in process, March 1 0 1116 1116
Cost of completed March 1 work in process 17298
Started and completed in March 208260 96120 304380
Transferred to Packing Department in March 321678
Inventory in process, March 31 9360 2592 11952
Total costs assigned by the Baking Department 333630
a
1. Direct materials cost per equivalent unit $1.30
2. Conversion cost per equivalent unit $0.60
3. Cost of the beginning work in process completed during March $17298
4. Cost of units started and completed during March $304380
5. Cost of the ending work in process $11952
b
Conversion cost in beginning inventory=16182-(9300*1.30)= $4092
Conversion cost per equivalent unit = 4092/(9300*4/5)= $0.55
The conversion cost per equivalent unit increase